OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 8) and a directrix at y = -8

OpenStudy (whpalmer4):



OpenStudy (whpalmer4):



OpenStudy (anonymous):

would the answer be y = 1/8x^2?

OpenStudy (whpalmer4):

The distance between the focus and the vertex, and the distance between the vertex and the directrix, is a constant \(p\). \[4p(y-k) = (x-h)^2\]where the vertex is at \((h,k)\)

OpenStudy (anonymous):

So 8 right?

OpenStudy (whpalmer4):

You tell me! But remember that the definition of a parabola is all the points equidistance from the focus and the directrix. Here's a graph of your parabola, and I've marked in two lines at x=8, y=8. The directrix is the purple line across the bottom. Does the point where the two lines intersect look like it is an equal distance from (0,8)?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

NO! From (0,8) to (8,8) is 8 units. From (8,-8) to (8,8) is 16 units.

OpenStudy (whpalmer4):

Why don't you show me how you came up with your formula, so we can figure out the mistake?

OpenStudy (anonymous):

Oh, I misread your comment. When you said " Does the point where the two lines intersect look equal" I looked at (0,8) to (8,8) and (0,8) to (8,-8)

OpenStudy (whpalmer4):

Oh, yeah, then it would look equal, I agree :-)

OpenStudy (whpalmer4):

But do you agree that the point over at (8,8) is not equally far from (0,8) and the nearest approach of the directrix at (8,-8)?

OpenStudy (anonymous):

agreed

OpenStudy (whpalmer4):

That means that point is not on the parabola we want.

OpenStudy (whpalmer4):

Do you know the vertex form equation for a parabola?

OpenStudy (anonymous):

y= a(x-h)^2 + k?

OpenStudy (whpalmer4):

That's one version of it, yes. We will find it more convenient to rearrange that a bit: \[y-k = a(x-h)^2\]\[\frac{1}{a}(y-k) = (x-h)^2\] We can still "read out" the vertex from that, agreed?

OpenStudy (anonymous):

yeah

OpenStudy (whpalmer4):

Okay, I'm going to change it to \[4p(y-k) = (x-h)^2\]\[4p=\frac{1}{a}\]\[p=\frac{1}{4a}\] \(p\) is the distance between the focus and the vertex, and the distance between the vertex and the directrix. We know those distances here, don't we?

OpenStudy (anonymous):

1/32?

OpenStudy (whpalmer4):

what's 1/32?

OpenStudy (anonymous):

p =1/4(8) = 1/32 So the answer would be y= 1/32 x^2?

OpenStudy (anonymous):

or am I jumping the gun?

OpenStudy (whpalmer4):

yes, \[y = \frac{1}{32}x^2\]is the correct answer.

OpenStudy (anonymous):

thank you

OpenStudy (whpalmer4):

In this graph, the blue line is your original answer, the purple is the new, improved answer, the olive line is the directrix, and the two black lines I added intersect at (16,8), which you can this time is equally far from the focus (0,8) and the directrix y = -8

OpenStudy (whpalmer4):

you got the right answer but did the math wrong :-) p = 8 here. p is the distance between focus and vertex: (0,8) to (0,0) = 8 and also the distance between vertex and directrix: (0,0) to y = -8 = 8

OpenStudy (whpalmer4):

If you have your equation in the \[4p(y-k)=(x-h)^2\]variant of vertex form, the vertex is at \((h,k)\), the focus is at \((h,k+p)\) and the directrix is at \(y = k-p\)

OpenStudy (anonymous):

but either way I will still end up with y=1/32 x^2 right?

OpenStudy (whpalmer4):

Yes, \[y=\frac{1}{32}x^2\] is the correct answer here. But \(p\ne \frac{1}{32}\) as you wrote above. Perhaps you meant that \(p=8\) and just were a bit sloppy in how you wrote your work.